# Astrometric properties of reference frame sources as a function of redshift

### v1

@@ Abstract
Methods. We compiled redshifts for about 10 000 RFC sources over the range 0<z<50\!<\!z\!<\!5 by combining data from the Dark Energy Spectroscopic Instrument Data Release 1 and the Sloan Digital Sky Survey Data Release 17/19 with additional datasets from the NASA/IPAC Extragalactic Database. Cross-matching with Gaia Data Release 3 yielded a clean sample of 4 068 RFC objects with reliable spectroscopic redshifts and classifications (galaxies and quasi-stellar objects; QSOs). We analysed the redshift dependence of their radio astrometric properties from very long baseline interferometry (VLBI) and their optical astrometric properties from Gaia.
@@ 1 Introduction
Recent studies have explored the connection between the physical properties, redshifts, and astrometric behaviour of reference sources (Makarov et al. 2019; Liu et al. 2021; Secrest 2022; Gaia Collaboration et al. 2023; Makarov 2024; Lambert and Secrest 2024; Lambert et al. 2025; Huda et al. 2025). In particular, Sexton et al. (2022) presented a detailed spectroscopic characterisation of a subset of ICRF3 sources based on Sloan Digital Sky Survey Data Release 16 (SDSS DR16; Ahumada et al. 2020). This work indicated the importance of reliable redshifts and spectral classifications for understanding the physical nature of reference frame objects and their astrometric behaviour. Extending such analyses to larger samples can improve our understanding of the redshift dependence of astrometric properties.
@@ 2.1 The radio fundamental catalogue
In this work, we adopt the release version rfc_2025c of the RFC catalogue [^1] https://astrogeo.org/sol/rfc/rfc_2025c/ as the parent sample for redshift determination and cross-matching with Gaia DR3 data. This release contains 21 949 sources, whose absolute positions and correlated flux densities were determined from the analysis of virtually all publicly available VLBI sessions conducted over the past four decades. The catalogue achieves milliarcsecond-level astrometric accuracy by applying a novel technique that combines and calibrates heterogeneous VLBI datasets, and by rigorously evaluating systematic and random errors through decimation tests, cross-comparisons, and error modelling (Petrov and Kovalev 2025). Its high precision and wide sky coverage make it fundamental to a wide range of applications, from celestial reference frame definition to space geodesy, spacecraft navigation, AGN studies, and modern VLBI phase-referencing observations.
@@ 2.2 DESI/SDSS spectroscopic data
We also make use of spectroscopic data from the seventeenth data release of SDSS (SDSS DR17; Abdurro’uf et al. 2022), the final release of the SDSS-IV programme. SDSS DR17 provides one of the most comprehensive spectroscopic datasets available, obtained with the 2.5 m Sloan Foundation Telescope at Apache Point Observatory. Compared to earlier releases such as DR13, SDSS DR17 represents a major expansion and refinement of the spectroscopic dataset. The number of reliable spectra in the main QSO sample from the Extended Baryon Oscillation Spectroscopic Survey (eBOSS; Dawson et al. 2016) has increased from roughly 34 00034\,000 in DR13 to more than 430 000430\,000 in DR17. Similarly, the spectroscopic sample of LRGs and ELGs has expanded from about 47 00047\,000 to over 460 000460\,000 objects. DR17 also incorporates significant methodological improvements, including the use of PCA-based spectral fitting and extensive visual inspection in addition to the standard SDSS pipeline, resulting in more accurate classifications, particularly for quasars. As the final data release of SDSS-IV, DR17 thus provides a more complete and higher-quality spectroscopic dataset than previous releases, enabling more robust studies of galaxies, quasars, and the large-scale structure of the Universe.
In addition, we also make use of spectroscopic data from the latest SDSS DR19 (SDSS Collaboration et al. 2025), which is part of the SDSS-V program. As the first substantial public release of new data from the SDSS-V ecosystem, DR19 provides updates relative to the final SDSS-IV release. The DR19 data include early observations from Black Hole Mapper (BHM), which provide expanded spectroscopic coverage of quasars and AGN. The new data and improved measurements of DR19 complement the DR17 dataset, allowing for more robust analyses.
@@ 2.3 Sample selection
The redshift information for RFC sources is assembled in a hierarchical manner during the sample selection process. We adopt a sequential strategy that prioritises the most reliable spectroscopic measurements. First, we use spectroscopic redshifts from the DESI DR1 catalogue with quality flag ${\tt ZWARN}=0$, resulting in 3 403 reliable redshifts. For sources without reliable DESI redshifts, we adopt spectroscopic redshifts from SDSS DR17 and SDSS DR19 with ${\tt ZWARNING}=0$, which provide an additional 1 832 sources. In total, this sequential procedure provides 5 235 sources with reliable spectroscopic redshifts from DESI or SDSS. All these spectroscopic redshifts are required to satisfy z>0z\!>\!0 and have measurement uncertainties smaller than 0.01.
@@
For sources lacking spectroscopic measurements, we supplement the sample with 4,672 redshifts compiled from the literature. Most of these literature-based redshifts were compiled from NED by the RFC cross-match table.
@@
Following this procedure, we obtain redshift information for 9 907 RFC sources in the catalogue. More than half of these redshifts are provided by the latest DESI and SDSS spectroscopic observations. This combined dataset provides broad coverage over the redshift range 0<z<50\!<z\!<5, ensuring both high reliability for the majority of the sample and extensive completeness for statistical analysis.
@@
To assess the reliability of the adopted spectroscopic redshifts, we perform a cross-checking between independent surveys (Appendix B). Overall, we find that the majority of sources show good agreement in redshifts across different methods ($|\Delta z|\!<\!0.1$). However, we find that DESI spectroscopic redshifts in the range 1.6<z<1.71.6\!<z\!<1.7 show significant deviations from redshifts measured by other methods. According to DESI Collaboration et al. (2025), DESI spectra in the range 1.6<z<1.631.6\!<z\!<1.63 are constrained only by the [O ii] $\lambda\lambda 3726,3729$ doublet and are affected by significant sky background, whereas spectra in the range 1.63<z<1.71.63\!<z\!<1.7 have no major emission line coverage. As a result, the DESI spectroscopic redshift solutions in this range are susceptible to unphysical fits and show systematic offsets compared with other determinations.
@@
To minimise the impact of such unreliable measurements, we first exclude all DESI redshifts within the range 1.6<z<1.71.6\!<z\!<1.7 from our sample. For the remaining sources that have both SDSS and DESI spectroscopic measurements, we retain only those for which the two redshifts are consistent within $|\Delta z|\!<\!0.1$. The consistency criterion of $|\Delta z|\!<\!0.1$ is chosen conservatively to flag catastrophic failures rather than statistical uncertainties. Additionally, for sources with DESI redshifts in the range 1.6<z<1.71.6\!<z\!<1.7 but still satisfy the consistency criterion, we retain these objects and use the SDSS redshift for our analysis. These conservative cuts ensure that our sample used in the statistical analysis is both robust and internally consistent across surveys. After applying these selection criteria, 464 sources are excluded, leaving a sample of 4 771 objects ($\sim\!90\%$) with reliable spectroscopic redshifts from either DESI or SDSS. By subsequently cross-matching with Gaia DR3, we obtain 4 068 objects that have spectroscopic redshifts from either SDSS or DESI, including 934 ICRF3 sources, which constitute our final sample.
@@
We are particularly interested in whether the spectral classification of RFC sources affects the relationship between their astrometric properties and spectroscopic redshift. For this purpose, we use the spectral classifications provided by SDSS and DESI, resulting in 784 galaxies, 3 151 QSOs, and 133 sources that are classified as stars or have inconsistent classifications between SDSS and DESI. Based on this classification, we further examine the relationships between the astrometric parameters of RFC sources and redshift.
@@ 2.4 Redshift distributions
After applying the sample selection and spectral classification, we present a series of redshift distributions in Fig. 1. A complementarity is seen in the redshift distribution of RFC sources: SDSS spectroscopic redshifts are primarily concentrated at low redshifts, whereas DESI spectroscopic measurements extend the redshift coverage toward higher redshifts. A deficit of DESI sources appears at 1.6<z<1.71.6\!<z\!<1.7, reflecting known reliability issues in this redshift range. The corresponding excess of SDSS redshifts in this range arises because we use the SDSS redshift for the sources in this range whose DESI and SDSS redshifts satisfy the $|\Delta z|\!<\!0.1$ criterion in our analysis. The distribution of the same sample separated by spectral classification shows that galaxies dominate the low-redshift population, close to the local Universe (0<z<0.50\!<z\!<0.5), whereas QSOs populate a broad redshift range extending to z>4z\!>\!4, with their highest concentration occurring at 1.0<z<1.51.0\!<z\!<1.5.
@@ 3 Results
We systematically analyse the redshift dependence of the astrometric parameters of RFC sources obtained from radio VLBI observations and optical Gaia DR3 measurements (binned with $\Delta z\!=\!0.2$). All the data are smoothed using the Nadaraya–Watson estimator to better illustrate the trends of the astrometric parameters as a function of redshift.
Figure 2: Positional uncertainty derived from RFC VLBI observations (left) and Gaia optical measurements (right) as functions of redshift. For each panel, the orange data points indicate QSOs, the green data points indicate galaxies in the sample, and the grey data points indicate sources that are either unclassified or lack a reliable classification. The blue and red solid lines represents the median values of the galaxies and QSOs, respectively, whereas the purple line represents the median value of the total sample. In addition, to ensure the reliability of the results, median values are not shown for bins containing fewer than 20 data points.
@@ 3.1 Astrometric positional precision
We compare the median positional uncertainties for RFC and Gaia DR3. The positional uncertainties ($\sigma_{\texttt{p}}$) of VLBI and Gaia are both defined as the semi-major axis of the positional error ellipse and are calculated as (Lindegren et al. 2016):
@@
We further examine the positional uncertainties of different types of RFC sources and find that QSOs at z>0.5z\!>\!0.5 show no significant redshift dependence in either the radio or optical bands. Galaxies have much larger positional uncertainties than QSOs in both radio and optical bands, and their optical uncertainties decrease significantly with redshift within 0<z<0.50\!<z\!<0.5. This suggests that the negative gradient in the optical positional uncertainties of the total sample within this redshift range is most likely driven by the contribution from galaxies.
We also analyse the redshift dependence of the structure index and compactness of ICRF3 sources in our sample separately at the $S$, $X$, $K$, and $Q$ bands (Fey and Charlot 2000). These two parameters characterise the radio morphological properties of the sources and are obtained from the Bordeaux VLBI Image Database (BVID; Collioud and Charlot 2019) [^2] https://bvid.astrophy.u-bordeaux.fr/. Since only a small number of sources in our sample have observations at the $K$ and $Q$ bands, reliable results are obtained only for the $S$ and $X$ bands. At the $S$ band, the median structure index is about 2, while at the $X$ band it is about 3. The median compactness of the sample is about 0.6 in both the $S$ and $X$ bands. We find that the structure index of our sample shows no significant correlation with redshift in any band.
@@ 3.2 Radio-optical offset
We calculate absolute and relative offsets following the same procedures as in Mignard et al. (2016), where these offsets are referred to as the angular separations and normalised separations, respectively. As shown in Fig. 3, the median absolute and relative radio-optical offsets of the total sample decrease rapidly with increasing redshift over the range 0<z<0.50\!<z\!<0.5, followed by a slower decrease over 0.5<z<1.30.5\!<z\!<1.3. The median absolute offset reaches a minimum around $z\!\simeq\!1.3$ and exhibits a slight increase over the range 1.3<z<3.01.3\!<z\!<3.0 (from 0.7 mas at $z\!\simeq\!1.3$ to 1.2 mas at $z\!\simeq\!3.0$). However, the median relative offset remains nearly constant beyond z>1.3z\!>\!1.3. This overall trend of absolute and relative offsets is consistent with the results reported by Makarov et al. (2019) for ICRF3 sources. Additionally, this behaviour is similar to the trends observed in the optical positional uncertainties (Fig. 2).
@@ 3.3 Parallaxes and proper motions
As shown in Fig. 4, we investigate the relationships between parallax and redshift. We find that the parallax of the sample shows no clear dependence on redshift, with the median values in different bins concentrated between 0 and 0.1-0.1 mas. For QSOs, the median parallax remains around 0.02-0.02 mas over the redshift range 0.5<z<1.50.5\!<z\!<1.5, which likely reflects the parallax zero-point offset present in Gaia measurements (Lindegren et al. 2021a; Ding et al. 2024). In both the total sample and the QSO subsample, the parallax uncertainty decreases significantly with increasing redshift over the range 0<z<0.50\!<z\!<0.5, followed by a slower decrease over 0.5<z<10.5\!<z\!<1. For z>1z\!>\!1, the parallax uncertainty increases slowly with redshift, rising from about $0.2$ mas at $z\!\sim\!1$ to about $0.3$ mas at $z\!\sim\!3.5$. In addition, compared with QSOs at comparable redshifts, galaxies exhibit a more negative median parallax and larger parallax uncertainties.
We also investigate the relationships between proper motion and redshift (Fig. 5). We find a similar behaviour to that of the optical positional uncertainties: both the median proper motion and proper motion uncertainty decrease rapidly over 0<z<0.50\!<z\!<0.5, followed by a slower decrease over 0.5<z<10.5\!<z\!<1, and exhibit only a slight increase at z>1z\!>\!1. This trend is consistent with the changing source composition, in which galaxies dominate at low redshift and exhibit larger values that decrease clearly with increasing redshift, whereas QSOs dominate at high redshift (z>1z\!>\!1) and exhibit smaller values with only a weak increase.
Furthermore, we compute the error-normalised proper motion $\chi$ as in Makarov and Secrest (2022):

### v2

# arXiv:2602.19135v2[astro-ph.IM] 28 Feb 2026

@@ Abstract
Methods. We compiled redshifts for about 10 000 RFC sources over the range 0<z<50\!<\!z\!<\!5 by combining data from the Dark Energy Spectroscopic Instrument Data Release 1 and the Sloan Digital Sky Survey Data Release 17/19 with additional datasets from the NASA/IPAC Extragalactic Database. Cross-matching with Gaia Data Release 3 yielded a sample of 4 068 RFC objects with reliable spectroscopic redshifts and classifications, including galaxies and quasi-stellar objects (QSOs). We analysed the redshift dependence of their radio astrometric properties from very long baseline interferometry (VLBI) and their optical astrometric properties from Gaia.
@@ 1 Introduction
Recent studies have explored the connection between the physical properties, redshifts, and astrometric behaviour of reference sources (Makarov et al. 2019; Liu et al. 2021; Secrest 2022; Gaia Collaboration et al. 2023; Makarov 2024; Lambert and Secrest 2024; Lambert et al. 2025; Huda et al. 2025). In particular, Sexton et al. (2022) presented a detailed spectroscopic characterisation of a subset of ICRF3 sources based on Sloan Digital Sky Survey Data Release 16 (SDSS DR16; Ahumada et al. 2020). This work indicates the importance of reliable redshifts and spectral classifications for understanding the physical nature of reference frame objects and their astrometric behaviour. Extending such analyses to larger samples can improve our understanding of the redshift dependence of astrometric properties.
@@ 2.1 The Radio Fundamental Catalogue
In this work, we adopted the release version rfc_2025c of the RFC catalogue [^1] https://astrogeo.org/sol/rfc/rfc_2025c/ as the parent sample for redshift determination and cross-matching with Gaia DR3 data. This release contains 21 949 sources, whose absolute positions and correlated flux densities were determined from the analysis of virtually all publicly available VLBI sessions conducted over the past four decades. The catalogue achieves milliarcsecond-level astrometric accuracy by applying a novel technique that combines and calibrates heterogeneous VLBI datasets, and by rigorously evaluating systematic and random errors through decimation tests, cross-comparisons, and error modelling (Petrov and Kovalev 2025). Its high precision and wide sky coverage make it fundamental to a wide range of applications, from celestial reference frame definition to space geodesy, spacecraft navigation, AGN studies, and modern VLBI phase-referencing observations.
@@ 2.2 DESI/SDSS spectroscopic data
We also made use of spectroscopic data from the seventeenth data release of SDSS (SDSS DR17; Abdurro’uf et al. 2022), the final release of the SDSS-IV programme. SDSS DR17 provides one of the most comprehensive spectroscopic datasets available, obtained with the 2.5 m Sloan Foundation Telescope at Apache Point Observatory. Compared to earlier releases such as DR13, SDSS DR17 represents a major expansion and refinement of the spectroscopic dataset. The number of reliable spectra in the main QSO sample from the Extended Baryon Oscillation Spectroscopic Survey (eBOSS; Dawson et al. 2016) has increased from roughly 34 00034\,000 in DR13 to more than 430 000430\,000 in DR17. Similarly, the spectroscopic sample of LRGs and ELGs has expanded from about 47 00047\,000 to over 460 000460\,000 objects. DR17 also incorporates significant methodological improvements, including the use of PCA-based spectral fitting and extensive visual inspection in addition to the standard SDSS pipeline, resulting in more accurate classifications, particularly for quasars. As the final data release of SDSS-IV, DR17 thus provides a more complete and higher-quality spectroscopic dataset than previous releases, enabling more robust studies of galaxies, quasars, and the large-scale structure of the Universe.
In addition, we also made use of spectroscopic data from the latest SDSS DR19 (SDSS Collaboration et al. 2025), which is part of the SDSS-V program. As the first substantial public release of new data from the SDSS-V ecosystem, DR19 provides updates relative to the final SDSS-IV release. The DR19 data include early observations from the Black Hole Mapper (BHM), which provide expanded spectroscopic coverage of quasars and AGN. The new data and improved measurements of DR19 complement the DR17 dataset, allowing for more robust analyses.
@@ 2.3 Sample selection
The redshift information for RFC sources was assembled in a hierarchical manner during the sample selection process. We adopted a sequential strategy that prioritised the most reliable spectroscopic measurements. First, we used spectroscopic redshifts from the DESI DR1 catalogue with quality flag ${\tt ZWARN}=0$, resulting in 3 403 reliable redshifts. For sources without reliable DESI redshifts, we adopted spectroscopic redshifts from SDSS DR17 and SDSS DR19 with ${\tt ZWARNING}=0$, which provided an additional 1 832 sources. In total, this sequential procedure provided 5 235 sources with reliable spectroscopic redshifts from DESI or SDSS. All these spectroscopic redshifts were required to satisfy z>0z\!>\!0 and have measurement uncertainties smaller than 0.01.
@@
For sources lacking spectroscopic measurements, we supplemented the sample with 4,672 redshifts compiled from the literature. Most of these literature-based redshifts were compiled from NED by the RFC cross-match table.
@@
Following this procedure, we obtained redshift information for 9 907 RFC sources in the catalogue. More than half of these redshifts were provided by the latest DESI and SDSS spectroscopic observations. This combined dataset provided broad coverage over the redshift range 0<z<50\!<z\!<5, ensuring both high reliability for the majority of the sample and extensive completeness for statistical analysis.
@@
To assess the reliability of the adopted spectroscopic redshifts, we performed cross-checking between independent surveys (Appendix B). Overall, we found that the majority of sources show good agreement in redshifts across different methods ($|\Delta z|\!<\!0.1$). However, we found that DESI spectroscopic redshifts in the range 1.6<z<1.71.6\!<z\!<1.7 show significant deviations from redshifts measured by other methods. According to DESI Collaboration et al. (2025), DESI spectra in the range 1.6<z<1.631.6\!<z\!<1.63 are constrained only by the [O ii] $\lambda\lambda 3726,3729$ doublet and are affected by significant sky background, whereas spectra in the range 1.63<z<1.71.63\!<z\!<1.7 have no major emission line coverage. As a result, the DESI spectroscopic redshift solutions in this range are susceptible to unphysical fits and show systematic offsets compared with other determinations.
@@
To minimise the impact of such unreliable measurements, we first excluded all DESI redshifts within the range 1.6<z<1.71.6\!<z\!<1.7 from our sample. For the remaining sources that had both SDSS and DESI spectroscopic measurements, we retained only those for which the two redshifts were consistent within $|\Delta z|\!<\!0.1$. The consistency criterion of $|\Delta z|\!<\!0.1$ was chosen conservatively to flag catastrophic failures rather than statistical uncertainties. Additionally, for sources with DESI redshifts in the range 1.6<z<1.71.6\!<z\!<1.7 but still satisfied the consistency criterion, we retained these objects and used the SDSS redshift for our analysis. These conservative cuts ensured that our sample used in the statistical analysis was both robust and internally consistent across surveys. After applying these selection criteria, 464 sources were excluded, leaving a sample of 4 771 objects ($\sim\!90\%$) with reliable spectroscopic redshifts from either DESI or SDSS. By subsequently cross-matching with Gaia DR3, we obtained 4 068 objects that had spectroscopic redshifts from either SDSS or DESI, including 934 ICRF3 sources, which constitute our final sample.
@@
We are particularly interested in whether the spectral classification of RFC sources affects the relationship between their astrometric properties and spectroscopic redshift. For this purpose, we used the spectral classifications provided by SDSS and DESI, resulting in 784 galaxies, 3 151 QSOs, and 133 sources that were classified as stars or had inconsistent classifications between SDSS and DESI. Based on this classification, we further examined the relationships between the astrometric parameters of RFC sources and redshift.
@@ 2.4 Redshift distributions
After applying the sample selection and spectral classification, we present a series of redshift distributions in Fig. 1. A complementarity is seen in the redshift distribution of RFC sources: SDSS spectroscopic redshifts are primarily concentrated at low redshifts, whereas DESI spectroscopic measurements extend the redshift coverage toward higher redshifts. A deficit of DESI sources appears at 1.6<z<1.71.6\!<z\!<1.7, reflecting known reliability issues in this redshift range. The corresponding excess of SDSS redshifts in this range arises because we used the SDSS redshift for the sources in this range whose DESI and SDSS redshifts satisfy the $|\Delta z|\!<\!0.1$ criterion in our analysis. The distribution of the same sample separated by spectral classification shows that galaxies dominate the low-redshift population, close to the local Universe (0<z<0.50\!<z\!<0.5), whereas QSOs populate a broad redshift range extending to z>4z\!>\!4, with their highest concentration occurring at 1.0<z<1.51.0\!<z\!<1.5.
@@ 3 Results
We systematically analysed the redshift dependence of the astrometric parameters of RFC sources obtained from radio VLBI observations and optical Gaia DR3 measurements (binned with $\Delta z\!=\!0.2$). All the data were smoothed using the Nadaraya–Watson estimator to better illustrate the trends of the astrometric parameters as a function of redshift.
Figure 2: Positional uncertainty derived from RFC VLBI observations (left) and Gaia optical measurements (right) as functions of redshift. For each panel, the orange data points indicate QSOs, the green data points indicate galaxies in the sample, and the grey data points indicate sources that are either unclassified or lack a reliable classification. The blue and red solid lines represent the median values of the galaxies and QSOs, respectively, whereas the purple line represents the median value of the total sample. In addition, to ensure the reliability of the results, median values are not shown for bins containing fewer than 20 data points.
@@ 3.1 Astrometric positional precision
We compared the median positional uncertainties for RFC and Gaia DR3. The positional uncertainties ($\sigma_{\texttt{p}}$) of VLBI and Gaia are both defined as the semi-major axis of the positional error ellipse and are calculated as (Lindegren et al. 2016):
@@
We further examined the positional uncertainties of different types of RFC sources and find that QSOs at z>0.5z\!>\!0.5 show no significant redshift dependence in either the radio or optical bands. Galaxies have much larger positional uncertainties than QSOs in both radio and optical bands, and their optical uncertainties decrease significantly with redshift within 0<z<0.50\!<z\!<0.5. This suggests that the negative gradient in the optical positional uncertainties of the total sample within this redshift range is most likely driven by the contribution from galaxies.
We also analysed the redshift dependence of the structure index and compactness of ICRF3 sources in our sample separately at the $S$, $X$, $K$, and $Q$ bands (Fey and Charlot 2000). These two parameters characterise the radio morphological properties of the sources and are obtained from the Bordeaux VLBI Image Database (BVID; Collioud and Charlot 2019) [^2] https://bvid.astrophy.u-bordeaux.fr/. Since only a small number of sources in our sample have observations at the $K$ and $Q$ bands, reliable results are obtained only for the $S$ and $X$ bands. At the $S$ band, the median structure index is about 2, while at the $X$ band it is about 3. The median compactness of the sample is about 0.6 in both the $S$ and $X$ bands. We find that the structure index of our sample shows no significant correlation with redshift in any band.
@@ 3.2 Radio-optical offset
We calculated absolute and relative offsets following the same procedures as in Mignard et al. (2016), where these offsets are referred to as the angular and normalised separations, respectively. As shown in Fig. 3, the median absolute and relative radio-optical offsets of the total sample decrease rapidly with increasing redshift over the range 0<z<0.50\!<z\!<0.5, followed by a slower decrease over 0.5<z<1.30.5\!<z\!<1.3. The median absolute offset reaches a minimum around $z\!\simeq\!1.3$ and exhibits a slight increase over the range 1.3<z<3.01.3\!<z\!<3.0 (from 0.7 mas at $z\!\simeq\!1.3$ to 1.2 mas at $z\!\simeq\!3.0$). However, the median relative offset remains nearly constant beyond z>1.3z\!>\!1.3. This overall trend of absolute and relative offsets is consistent with the results reported by Makarov et al. (2019) for ICRF3 sources. Additionally, this behaviour is similar to the trends observed in the optical positional uncertainties (Fig. 2).
@@ 3.3 Parallaxes and proper motions
As shown in Fig. 4, we investigated the relationships between parallax and redshift. We find that the parallax of the sample shows no clear dependence on redshift, with the median values in different bins concentrated between 0 and 0.1-0.1 mas. For QSOs, the median parallax remains around 0.02-0.02 mas over the redshift range 0.5<z<1.50.5\!<z\!<1.5, which likely reflects the parallax zero-point offset present in Gaia measurements (Lindegren et al. 2021a; Ding et al. 2024). In both the total sample and the QSO subsample, the parallax uncertainty decreases significantly with increasing redshift over the range 0<z<0.50\!<z\!<0.5, followed by a slower decrease over 0.5<z<10.5\!<z\!<1. For z>1z\!>\!1, the parallax uncertainty increases slowly with redshift, rising from about $0.2$ mas at $z\!\sim\!1$ to about $0.3$ mas at $z\!\sim\!3.5$. In addition, compared with QSOs at comparable redshifts, galaxies exhibit a more negative median parallax and larger parallax uncertainties.
We also investigated the relationships between proper motion and redshift (Fig. 5). We find a similar behaviour to that of the optical positional uncertainties: both the median proper motion and its median uncertainty decrease rapidly over 0<z<0.50\!<z\!<0.5, followed by a slower decrease over 0.5<z<10.5\!<z\!<1, and exhibit only a slight increase at z>1z\!>\!1. This trend is consistent with the changing source composition, in which galaxies dominate at low redshift and exhibit larger values that decrease clearly with increasing redshift, whereas QSOs dominate at high redshift (z>1z\!>\!1) and exhibit smaller values with only a weak increase.
Furthermore, we computed the error-normalised proper motion $\chi$ as in Makarov and Secrest (2022):